# Firstname1 Lastname1    Firstname2 Lastname2    Firstname3 Lastname3    Firstname4 Lastname4    Firstname5 Lastname5    Firstname6 Lastname6    Firstname7 Lastname7    Firstname8 Lastname8    Firstname8 Lastname8

### v1

@@ 4.2 Interpretability: Fixed Modulus
Qualitative Algorithm (fixed modulus): i. Find mixed-radix representations of inputs from the learned prime factorization of $m$ ii. Look back $r=p_{j}^{k}$ steps in the context and copy the lowest $k$ digits, for different prime factors of $m$ iii. Using these $r$-step sequences, predict the higher digits of the simplified sequence
@@
We now discuss the algorithm implemented by Transformers trained on LCG with a fixed modulus. We will show that, despite being provided integers as inputs, the model develops emergent structures that create and leverage the mixed-radix representation of the inputs. We will focus on the setup with a 1-layer, single-head Transformer trained on LCG with $m=2048$. Similar results for composite moduli (e.g. $m=7776$) and different model sizes are presented in Section D.1.
@@
Figure 3: Test accuracy of trained model ($m=2048,l=1,n_{head}=1,d_{model}=768$), averaged over $a,c$ and initial seeds. (a) Test accuracy w.r.t. token positions. A ladder-like structure appears, with jumps occurring at $2^{k}$-th positions. (b) We represent numbers as an eleven-digit binary number ($2048=2^{11}$) and compute the per-digit test accuracy of model predictions.
@@
We begin by analyzing the average accuracy of a trained Transformer model as a function of token position along the context, shown in Figure 3 (a). The accuracy exhibits a ladder-like structure, with each jump occurring exactly at the $2^{k}$-th token position. These successive transitions can be explained using binary representations and $r$-step recurrence (Equation 4). Specifically, once the token position exceeds $2^{k}$, the model can look back in the context and implement $r=2^{k}$-step iteration. This is utilized for (i) copying the lowest $k$ bits, since they become constant as well as (ii) simplifying the higher bits, since their periods get reduced by a factor $2^{k}$. We note that the accuracy trend remains unchanged across different choices of $a$ and $x_{0}$ (see Section D.1 Figure 18 ).
@@
Next, we compute the per-digit accuracy by converting both model predictions and ground truth labels to their binary representations according Equation 2. In Figure 3(b), we observe that the model can predict more digits correctly with more context, with the transitions occurring at $2^{w}$-th token positions. This is a direct result of the sequences becoming increasingly simplified as the model can look farther back in the context and utilize $(r=2^{w})$-step iterations. Since this simplifications occurs in the form of digit-wise periods, Figure 3(b) serves as direct evidence that the model is internally developing and utilizing binary representations. The sequential learning of digits also explains the ladder-like structure of the accuracy in Figure 3(a). We find that the overall average accuracy (Figure 3(a)) multiplicatively depends on the per-digit accuracies (Figure 3(b)) (see Section D.1 Figure 15(b)) [^1] This holds if the per-digit accuracies are independent, which we confirm empirically.
@@
Figure 4: Fixed modulus interpretability ($m=2048$, depth=1, $n_{heads}=1,d_{model}=768,a=1589,c=629$). (a1) 1st principal component groups the numbers mod 2. (a2) 2nd and 3rd principal components group the numbers mod 4. (b) Embedding vectors of different numbers exhibit high cosine similarity when they are spaced $2^{k}$ distance apart – with the similarity increasing with $k$. (c) Each query attends most strongly to the tokens $2^{k}$ and $2^{k-1}$ distance backward, for highest possible value of $k$, enabling copying of $k$ lower bits. (d) post-ReLU MLP activations at token position 129 with respect to different target numbers (i.e. number at the next token position). Each neuron gets activated only for specific target numbers, with a preiodic pattern.
@@
Step i: We begin by analyzing how the model performs prime factorization to develop the binary representations (mixed-radix for general $m$). We conduct Principal Components Analysis (PCA) of the embedding matrix. Figure 4(a,b) shows the projections of all numbers $n\in\{0,\dots 2047\}$ along the first principle component of the embedding. We observe that the model groups the numbers into modulo $2$ clusters along the first principal component. Similarly, the 2nd and 3rd principal components group the numbers into modulo $4$ clusters. In general, we find principal directions that group the numbers into modulo $2^{w}$ clusters (see Section D.1 Figure 19). By clustering the numbers according to their remainder modulo different prime-powers, these principle directions naturally encode the digit-wise representations of the inputs. In Figure 4(c), we check the cosine similarity between the embedding vectors of different numbers. We see that the more digits two numbers share in the binary representation, the higer the cosine similarity of their embedding. This is a consequence of these numbers having similar components along principle directions corresponding to those digits.
Step ii, iii: To validate that the model looks back $r=2^{k}$ steps in the context, we examine the attention weights in Figure 4(d). In order to predict the token at position $t$, the model attends to the positions $t-{2^{k}}$. This allows the model to copy the lowest $k$ bits and simplify the prediction of higher bits. We observe that the model attends most strongly to the positions $t-{2^{k}}$ and $t-{2^{k-1}}$ for the largest possible value of $k$. This corresponds to the two brightest lines in Figure 4(d) that are ${2^{k-1}}$ distance apart. By comparing the binary representations of these two numbers, the model copies the lower $k$ bits to make predictions.
In addition to the two bright lines in the attention score, we observe multiple faint lines which are $2^{k^{\prime}}$ apart, for $k^{\prime}<k-1$. The model combines the information from all these tokens to predict the higher bits. To verify that the two bright lines at positions $t-{2^{w}}$ and $t-{2^{w-1}}$ copy lower bits, and the other faint lines are important for predicting higer bits, we perform the following ablation. For each query, we mask out all attention scores except for the top two key positions. We then measure the performance of this ablated model in Section D.1, Figure 15(d). In this case, we observe that the model retains the ability to copy the $k$ lowest bits, but loses the ability to predict higher bits.
After the attention layer gathers the information about previous tokens, MLPs implement the remaining computational processing to make predictions. In Figure 4 (d), we plot the post-ReLU MLP activations at fixed token positions w.r.t. the number appearing at that position. We find that each neuron only fires when a specific set of numbers appear at those tokens – often exhibiting periodic structure. We conclude that the first fully connected layer of the MLP determines which neurons to fire based on the context. The second fully connected layer aggregates the non-zero activations to form the final prediction.
@@ 4.3 Interpretability: Generalization to Unseen Modulus
Qualitative Algorithm (unseen modulus): i. Encode information about many possible prime factorizations ii. Use the largest number in the context to estimate the modulus iii. Implement steps ii and iii from the fixed modulus algorithm with a different realization
@@
Unlike the FM case, the training set for UM is generated using many different moduli $m$. Since each modulus has its own mixed-radix representation, which is incompatible with the other moduli, the model needs to implement a different, more general algorithm to solve UM tasks.
@@
We analyze how various components of a 4-layer Transformer model implement the above algorithm to solve the UM task. The model is trained on a dataset generated with $n_{m}=n_{a}=n_{c}=128$, where $m$ values are randomly chosen from the interval $[257,2457]$. To avoid leakage between train and test sets, we specifically exclude the moduli $m_{\mathrm{test}}\in\{1800=2^{3}\cdot 3^{2}\cdot 5^{2},2048=2^{11},2352=2^{4}\cdot 3\cdot 7^{2}\}$ from the training set.
((a))
Figure 5: PCA analysis of the embedding layer. The first principal component resembles a circular structure, while the second and third dimensions categorize numbers based on their modulo $2$ and modulo $3$ values, respectively.
@@
Step i: We begin by analyzing the embedding layer. In Figure 5(a), PCA reveals that the model forms a semi-circular structure in the first two principal components. This structure resembles the circular patterns observed in modular arithmetic tasks (power2022grokking; zhong2023clock). However, since our model is trained with multiple moduli, it cannot determine the value that should be identified with 00 to form a closed circle.
Further analysis of Figure 5 (a, b) demonstrates that the $2$nd and $3$rd principal components of the embedding layer categorize numbers according to their values modulo $2$ and $3$, respectively. Given that the modulus is not fixed in this case, there is no inherent requirement for the model to encode numbers in binary representation. In fact, as we will demonstrate next, the model forms radix representations based on the choice of modulus $m_{\mathrm{test}}$ through specialized attention heads. We believe the embedding layer’s primary focus on modulo $2$ and $3$ can be attributed to these being the most common prime factors of numbers appearing in the training set.
In the first layer, attention heads further process the embedded numbers by grouping them according to their modular values with respect to various prime factors, with each head specializing in specific prime factors. This head specialization enables the model to construct radix representations requiring different prime bases.
((a))
@@
Figure 6: PCA analysis of attention heads specializing in different prime factorizations and their effect on per-digit accuracy. (a1, b1, c1) PCA patterns of outputs extracted from corresponding heads. The first two principal components group numbers based on their values modulo $14$, modulo $4$, and modulo $3$, respectively. (a2, b2, c2) Test accuracy for individual digits when numbers are represented according to Equation 3 for given $m_{\mathrm{test}}$. (a3, b3, c3) Test accuracy for individual digits after pruning the corresponding heads shown in (a1, b1, c1). Pruning these specific heads substantially degrades performance on their corresponding digits.
The attention heads responsible for preparing different prime factorizations are shown in Figure 6. To get the (a1, a2, a3) panel, we feed the model with sequences generated using the corresponding $m_{\mathrm{test}}$ value and randomly selected legitimate values of $a$ and $c$. Then, we extract outputs from each corresponding head, at the identical token positions for all $m_{\mathrm{test}}$ possible numbers to form a $m_{\mathrm{test}}\times d_{\mathrm{model}}$ matrix. Finally, we perform PCA analysis over this matrix and project each number’s feature vector onto the top two principal components. The analysis reveals that each head naturally categorizes numbers according to their values modulo different prime factors. This head specialization enables the subsequent model layers to select which representation to use. The prominence of small prime numbers in the top principal components likely reflects their frequent occurrence in the prime factorization of composite numbers.
To further confirm the importance of these specialized heads, we compare the model’s per-digit performance before and after pruning the corresponding heads. In Figure 6 (a2, b2, c2), we measured the model’s performance before pruning, where we measured per-digit performance by representing numbers in the natural radix representation for each test modulus. The results show that the model typically struggles with calculating higher digits for a larger prime factor. This behavior aligns with what we observed in the FM case, where computing higher digits is more demanding than copying lower digits. Different from FM case, the difficulty in calculating higher digits is partly due to the model’s inability to precisely determine the modulus $m_{\mathrm{test}}$ from the sequence, which we will elaborate more later in this section.
In Figure 6 (a3, b3, c3), we measured the model’s performance after pruning the corresponding heads. By pruning, we mean replacing all components of the outputs from the corresponding heads with a single value. This value represents the mean magnitude of the outputs, computed by averaging across all dimensions: different sequences (using $10\%$ of randomly selected training sequences), token positions, and embedding dimension. Note that zeroing out any head’s contribution in the first layer would annihilate the model’s performance. In contrast, our pruning method ensures that we remove the contribution of an attention head without changing the scale of the signal drastically. Comparison with the pre-pruning results reveals that for each $m_{\mathrm{test}}$, pruning the head responsible for a specific prime factorization dramatically reduces performance on the corresponding digits while having a relatively minor impact on others. For instance, in panel (a3), the model’s ability to compute digits related to base-$7$ is completely lost, while digits related to base-$2$ and base-$3$, though impacted, remain above random chance. Similarly, panel (b3) shows complete performance degradation. In panel (c3), the model’s ability to compute digits related to $3$ is drastically decreased, while its capacity to process digits for base-$2$ and base-$5$ digits remains intact or shows marginal degradation.
To further validate the correlation between specific attention heads and digit-wise performance, we conduct pruning experiments as shown in Figures 20 and 21. Interestingly, in Figure 20 (a, b), when we remove the attention head responsible for computing modulo $3$ representations in Figure 6 (c1) and evaluate the model’s performance for $m_{\mathrm{test}}=2048$, we observe improved performance at specific token positions for certain bits, while other bits exhibit minimal performance degradation. This behavior suggests that the model experiences internal uncertainty about which radix representation to employ, likely stemming from its difficulty in accurately determining the value of $m_{\mathrm{test}}$, as we will demonstrate in the following paragraph.
((a))
@@
Figure 7: Attention head (layer 1, head 6) for estimating $m_{\mathrm{test}}$. (a) The query consistently attends to keys corresponding to the largest values within the given sequence; (b) the head generates features with the highest cosine similarity to $m_{\mathrm{test}}$; (c) when patching the feature with sequences generated using $m_{\mathrm{patch}}=1024$ but maintaining the same $a$ and $c$, the model often makes predictions with values of $n$ smaller than $m_{\mathrm{patch}}$.
Step ii: Next, we show how the model estimates $m_{\mathrm{test}}$ in-context in Figure 7, where we find one specific attention head in the first layer that is responsible for estimating $m_{\mathrm{test}}$. In panel (a), we visualize the attention weights for an input sequence generated with $a=5$, $c=31$, and $m_{\mathrm{test}}=2048$. We find that the head focuses exclusively on the few largest numbers within the context. In panel (b), using the same sequence, we extract output features from this specific head by zeroing out the contributions from all other heads in the first layer, and compute the cosine similarity between this extracted feature and the embedding of all numbers $n$. The head consistently maintains high cosine similarity with the few largest numbers observed within the context, representing a greedy estimation of $m_{\mathrm{test}}$.
@@
To further confirm that this head generates features corresponding to $m_{\mathrm{test}}$, we perform a patching experiment (zhang2024actpatch), as shown in panel Figure 7 (c). Here, we generate a sequence with the same $a$ and $c$ but with a different modulus $m_{\mathrm{patch}}=1024$, then use the feature from this patching sequence to replace the feature from the same head in the forward pass of the original sequence. The results show that the model frequently predicts subsequent numbers with values smaller than $m_{\mathrm{patch}}$. We also show in Figure 22 that patching any other heads might ruin the predictions but would never lead to a similar qualitative change, which further confirms that this head is the only one that helps the model estimate $m_{\mathrm{test}}$.
Step iii: After preparing the necessary features, subsequent layers in the model implement the rest of the algorithm. As in the FM case, we observe ladder structures in digit-wise accuracy for the lower digits at early token positions. This pattern, which demonstrates the copying bias, is visible in Figure 6 (b2). However, higher digits do not exhibit such ladder structures as the FM case. This discrepancy may arise because the model cannot precisely determine $m_{\mathrm{test}}$, requiring it to dynamically select the radix representation by identifying the most coherent signal at later stages. The potential competition mentioned earlier in step i further supports this argument. Nevertheless, we believe the model implements a similar algorithm for computing the higher bits in this case, which we will elaborate on in Section D.2.

### v2

@@ 4.2 Interpretability: Fixed Modulus
Qualitative Algorithm (fixed modulus): i. Find RNS representations of inputs from the learned prime factorization of $m$ ii. Look back $r=p_{j}^{k}$ steps in the context and copy the lowest $k$ digits, for different prime factors ($p_{j}$) of $m$ iii. Using these $r$-step iterations, predict the higher digits of the simplified sequence
@@
We now discuss the algorithm implemented by Transformers trained on LCG with a fixed modulus. We will show that, despite being provided integers as inputs, the model develops emergent structures that create and leverage the RNS representation of the inputs. We will focus on the setup with a 1-layer, single-head Transformer trained on LCG with $m=2048$. Similar results for composite moduli (e.g. $m=7776$) and different model sizes are presented in Appendix G. We emphasize that this algorithm works for arbitrary $a,c,x_{0}$ for a given $m$.
@@
Figure 4: FM: Test accuracy for $m=2048$, $\mathrm{depth}=1$, $n_{\mathrm{heads}}=1$, $d_{\mathrm{model}}=768$, averaged over $a$, $c$, and seeds. (a) Test accuracy w.r.t. token positions. Ladder-like structure appears, with jumps occurring at $2^{k}$-th positions. (b) We represent numbers as an eleven-digit binary number ($2048=2^{11}$) and compute the per-digit test accuracy of model predictions.
@@
We begin by analyzing the average accuracy of a trained Transformer model as a function of token position along the context, shown in Figure 4 (a). (Recall that in this section a token corresponds to an integer of the LCG sequence). The accuracy exhibits a ladder-like structure, with each jump occurring exactly at the $2^{k}$-th token position. These successive transitions can be explained using binary representations and $r$-step recurrence (Equation 3). Specifically, to predict the token at position $t\geq 2^{k}$, the model can look back in the context at position $t-2^{k}$ and implement $r=2^{k}$-step iteration. This allows the model to (i) copy the lowest $k$ bits since they remain unchanged; and (ii) simplify the higher bits, since their periods get reduced by a factor $2^{k}$. We note that the accuracy trend remains unchanged across different choices of $a$ and $x_{0}$ (see Figure 22).
@@
Next, in Figure 4(b), we compute the per-digit accuracy by converting both model predictions and ground truth labels to their binary representations according to Equation 2. We observe that the model can predict more digits correctly with more context, with sharp transitions occurring at $2^{k}$-th token positions. This is a direct result of the sequences becoming increasingly more simplified as the model can look farther back in the context and utilize $(r=2^{k})$-step iterations. Since these simplifications occur in the form of digit-wise periods, Figure 4(b) serves as direct evidence that the model is internally developing and utilizing binary representations. The sequential learning of digits also explains the ladder-like structure of the accuracy in Figure 4(a). We find that the overall average accuracy (Figure 4(a)) multiplicatively depends on the per-digit accuracies (Figure 4(b)) (empirical proof in Figure 23(b))  [^1] This holds if the per-digit accuracies are independent, which we confirm empirically.
@@
Figure 5: FM: ($a=1589,c=629$) Embedding layer. (a1) 1st principal component groups the numbers mod 2. (a2) 2nd and 3rd principal components group numbers mod 4. (b) Embedding vectors of different numbers exhibit high cosine similarity when they are spaced $2^{k}$ apart, with the similarity increasing with $k$.
@@
Step i: We begin by conducting Principal Components Analysis (PCA) of the embedding matrix, which shows how the model performs prime factorization to develop the binary representations (RNS for general $m$). Figure 5(a1) shows the projections of all numbers $x\in\{0,\dots 2047\}$ along the first principal component of the embedding. We observe that the model groups the numbers into modulo $2$ clusters along the first principal component. Similarly, the 2nd and 3rd principal components group the numbers into modulo $4$ clusters (Figure 5(a2)). In general, we find principal directions that group the numbers into modulo $2^{(\cdot)}$ (see Figure 25). By clustering the numbers according to their remainder modulo different prime-powers, these principal directions naturally encode the digit-wise representations of the inputs. In Figure 5(b), we check the cosine similarity between the embedding vectors of different numbers. We see that the more digits two numbers share in the binary representation, the higher the cosine similarity of their embedding. This is a consequence of these numbers having similar components along principal directions corresponding to those digits. For composite moduli, we find similar clustering according to different prime factors of $m$ (see Figures 26, 27).
Step ii, iii: In Figure 7(a) we examine the attention weights and find that to predict the number at position $t$, the model attends most strongly to the position $t-2^{k}$ for the highest possible value of $k$ s.t. $t\geq 2^{k}$ (i.e. $k=\lfloor\log_{2}{t}\rfloor$). This corresponds to the brightest line in Figure 7(a). Using the binary representation of the $(t-2^{k})$-th token, the model can copy the lowest $k$ bits and simplify the prediction of higher bits. Additionally, the second brightest line appeares at the position $t-2^{k-1}$, along with other faint lines at intermediate distances (multiples of $t-2^{k’}$ for $k’<k-1$). The information obtained from all these lines are utilized by the model in predicting the higher bits.
To verify that the brightestes line enables the copying of lower bits and that the second brightest line facilitates the prediction of higher bits, we performed the following two experiments on the trained model. (i) For each query in the attention head, we mask (zero) out the attention weights to all keys except for the one at position $t-2^{k}$. We then measure the performance of this ablated model, shown in Figure 6(a). We observe that the model retains the ability to copy the $k$ lowest bits, but loses the ability to predict the higher bits. (ii) Next, we repeat the above ablation experiment while masking out the attention to all keys except the two positions $t-2^{k}$ and $t-2^{k-1}$, shown in Figure 6(b). This results in a drastic improvement in the prediction of the higher bits compared to (i), confirming our assertions.
(a) Only attend to position $t-2^{k}$, where $k=\lfloor\log_{2}t\rfloor$ (brightest line). The model can copy the last $k$ bits, but cannot predict higher bits.
(b) Only attend to positions $t-2^{k}$ and $t-2^{k-1}$, where $k=\lfloor\log_{2}t\rfloor$ (top 2 brightest lines). Notably, the model can predict higher bits with good accuracy.
Figure 6: Per-digit test accuracies with attention masking, same model as Figure 4 ($m=2048$). For each query, attention to only some selected keys is kept. All other attention weights are masked.
After the attention layer collects information about previous tokens, the MLP block  [^2] MLP block: fully connected $\rightarrow$ ReLU $\rightarrow$ fully connected. processes the information to make predictions. We find that each hidden neuron (post-ReLU) in the MLP exhibits a periodic response with a distinct period, as a function of the target prediction. Consider the MLP at token position $t$. The input to the transformer at this position is $x_{t}$ and the output $y_{t}$, should match the next number in the sequence $x_{t+1}$. In Figure 7(b) we show the activation value of selected neurons at token position $t=129$ as a function of the target $x_{130}$, for $m$ different sequences obtained by changing the seed $x_{0}$ for a given LCG sequence. We observe that only a sparse set of neurons are activated for a given target $x_{130}$ and that there is a strong spiked periodic structure in the response of each neuron as a function of $x_{130}$, with neuron-specific frequencies.
The subsequent fully connected layer in the MLP block aggregates the contributions from all the activated neurons to make the correct prediction (McCracken et al., 2025). To visualize the contributions from an individual neuron, we mask out the contribution from all other neurons and extract the MLP output for a fixed input sequence. We then project this output onto the (un)embedding matrix, which shows us the contribution of that single neuron in making the correct prediction. [^3] We read-out MLP outputs instead of network outputs to avoid distortion from the skip connection. In Figure 7(c) we show that each neuron resolves the target number $x_{130}$ up to a distinct periodic pattern. The periodic patterns from different neurons constructively interfere at the target. In Figure 7(d), we observe that gradually adding contributions from multiple neurons resolves the correct output with increasing accuracy.
Figure 7: FM: ($m=2^{11},a=1589,c=629$) (a) Attention weights: each query attends most strongly to the tokens $2^{k}$ and $2^{k-1}$ distance backward, for the highest possible value of $k$, enabling copying of lowest $k$ bits. The other faint lines facilitate the prediction of higher bits. (b) Post-ReLU hidden layer MLP activations at token position $t=129$ (extracted using sequences with different $x_{0}$) as a function of the target number $x_{130}$ which it is supposed to predict. Each neuron gets activated only while predicting a specific $x_{130}$, exhibiting a sparse, periodic pattern. (c) Output of the MLP block projected onto the (un)embedding matrix; after masking out all but a single given hidden-layer neuron. The green dot denotes the value at the target number. Each neuron resolves the correct prediction up to a periodic structure. (d) Output of the MLP block projected onto the (un)embedding matrix; after combining the signal from multiple neurons (i.e. gradually un-masking the neurons). The per-neuron periodic patterns constructively interfere at the correct output.
@@ 4.3 Interpretability: Generalization to Unseen Modulus
Qualitative Algorithm (unseen modulus): i. Encode information about various possible prime factorizations ii. Estimate the modulus via the largest number in context iii. Combine steps i and ii to construct correct RNS representations, then implement steps ii and iii from the fixed modulus algorithm
@@
Unlike the FM case, the training set for UM is generated using many different moduli $m$, unseen at test time. Since each modulus has its own RNS representation incompatible with other moduli, the model must implement a different, more general algorithm to solve UM tasks.
@@
We analyze how the embedding layer and attention heads of a 4-layer Transformer model help implement the above algorithm to solve the UM task. The model is trained on a dataset generated with $n_{m}=n_{a}=n_{c}=128$, where $m_{\max}=2457$. To avoid leakage between training and test sets, we specifically exclude the moduli $m_{\mathrm{test}}\in\{1800=2^{3}\cdot 3^{2}\cdot 5^{2},2048=2^{11},2352=2^{4}\cdot 3\cdot 7^{2}\}$ from the training set. Note that the choice $m_{\mathrm{max}}=\lfloor 1.2\cdot 2048\rfloor$ is made to maintain consistency with the setting in Figure 1(a).
Figure 8: UM: PCA analysis of the embedding layer.
@@
Step i: We first analyze the embedding layer. In Figure 8(a), PCA shows a semi-circular structure along the first principal component, with the second principal component separating even and odd numbers. This semi-circle resembles the circular patterns seen in modular arithmetic tasks (Power et al., 2022; Zhong et al., 2023), but since our model is trained on multiple moduli, it cannot form a closed circle by identifying a unique 00 value. Figure 8(b) further shows that the $2$nd and $3$rd principal components group numbers by their remainders modulo $2$ and $3$, which likely reflects their prevalence as prime factors in the training set.
Note with varying moduli, the model need not form a binary encoding as before. Instead, we find attention heads in first layer group embedded numbers by remainders modulo different primes, each head specializing in a particular factor. This specialization allows the model to construct RNS representations with various prime bases.
To investigate head specialization, we analyze afformentioned first-layer attention heads in Figure 9, presenting PCA results for selected heads in panels (a1, b1, c1). For each, we input sequences using the corresponding $m_{\mathrm{test}}$ and randomly selected $(a,c)$ pairs (per the Hull-Dobell Theorem), then isolate the output $\bm{H}^{(h)}$ for each head by zeroing out all other heads. We perform PCA at token position $t=0$ ($\texttt{PCA}(\bm{H}^{(h)}[:,0,:])$), then projecting each number’s feature vector $\bm{H}^{(h)}[x,\,t,\,:]$ onto the top two principal components and labeling each point with its corresponding $x$.
The analysis shows that each head groups numbers by their remainder modulo different prime factors, enabling the model to select suitable representations in later layers. The prominence of small primes in the top principal components likely reflects their frequency in the training data. Furthermore, we believe the performance gains observed in Figure 1 with more heads can be partly attributed to the model’s improved capacity to capture additional prime factors for constructing RNS representations. More examples of specialization across various $(a,c,m_{\mathrm{test}},x_{0},t)$ are provided in Figure 28 (Section H.1).
@@
Figure 9: UM: PCA analysis of attention heads specialized for different prime factors and their impact on per-digit accuracy. (a1, b1, c1) PCA of outputs from specific heads; the first two principal components group numbers by remainder modulo $14$, $4$, and $3$, respectively. (a2, b2, c2) Test accuracy for individual digits with representations from Equation 4 for each $m_{\mathrm{test}}$. (a3, b3, c3) Per-digit test accuracy after pruning these heads, showing substantial performance degradation on the affected digits.
To further demonstrate that these heads directly influence the model’s performance, we measured per-digit accuracy before and after pruning each specialized head, as shown in panels (a2, b2, c2) and (a3, b3, c3). For pruning, we replace the head’s output with its mean value $\texttt{mean}(\bm{H}^{(h)}[:,\,:,\,:])\in\mathbb{R}$ across all positions (using $10\%$ of randomly selected training sequences), which preserves signal scale and avoids catastrophic model degradation. The results show that pruning a head responsible for a particular prime factor significantly impairs performance on corresponding digits, while other digits are less affected. For instance, in panel (a3), the model’s ability to compute $7^{1}$ and $7^{2}$ digits is lost, while base-2 and base-3 digits remain above chance. Similarly, in (b3), removing the modulo-2 head erases all corresponding digit accuracy, and in (c3), pruning the modulo-3 head greatly reduces base-3 digit performance, but leaves base-2 and base-5 largely intact.
Finally, we further validate the link between head specialization and digit-wise accuracy by pruning irrelevant heads. For $m_{\mathrm{test}}=2048$, pruning heads responsible for modulo $7$ or $3$ (as in panels (a1) or (c1)) yields the results in Figure 29. Interestingly, removing the modulo-3 head sometimes improves performance for specific token positions, while others show minimal degradation. These findings reinforce the connection between head specialization and digit-wise computation; further details can be found in Section H.2.
@@
Figure 10: UM: Attention head (layer 1, head 6) specialized in estimating $m_{\mathrm{test}}$. (a) Queries attend to largest keys; (b) the head produces features with the highest cosine similarity to $m_{\mathrm{test}}$; (c) patching this head with features from sequences with $m_{\mathrm{patch}}=1024$ (same $a$, $c$) steers the model to predict numbers $n<m_{\mathrm{patch}}$.
Step ii: The performance variations observed above suggest that the model is internally uncertain about which RNS representation to use, likely due to difficulty in determining $m_{\mathrm{test}}$. As we discuss below, the model appears to estimate $m_{\mathrm{test}}$ greedily via in-context learning.
In Figure 10(a), using a sequence with $x_{0}=1$, $a=5$, $c=31$, and $m_{\mathrm{test}}=2048$, we observe a first-layer attention head that attends primarily to the largest numbers, as seen by the vertical lines in the attention weights. To analyze this further, in panel (b), we extract the output $\bm{H}^{(h)}\in\mathbb{R}^{1\times L\times d_{\mathrm{model}}}$ from this head for the same sequence and compute its cosine similarity with token embeddings for all $x<m_{\mathrm{test}}$, producing a $\mathbb{R}^{L\times m_{\mathrm{test}}}$ matrix for the heatmap. This reveals that the head’s output consistently has the highest similarity with the largest numbers that has been seen in context, indicating a greedy estimation of $m_{\mathrm{test}}$.
@@
To further verify that this head estimates $m_{\mathrm{test}}$ for later layers, we conduct a patching experiment (Zhang & Nanda, 2024), as shown in Figure 10(c). We generate a new sequence with the same $a$, $c$, and $x_{0}$, but with $m_{\mathrm{patch}}=1024$, and extract $\bm{H}^{(h)}{\mathrm{patch}}\in\mathbb{R}^{1\times L\times d{\mathrm{model}}}$ from this head. We then overwrite the output $\bm{H}^{(h)}$ in a forward pass for the original sequence ($m_{\mathrm{test}}=2048$) with $\bm{H}^{(h)}{\mathrm{patch}}$. The model now frequently predicts numbers smaller than $m{\mathrm{patch}}$. As shown in Figure 31 (Section H.3), patching other heads disrupts predictions but never induces a similar qualitative shift, confirming this head’s unique role in estimating $m_{\mathrm{test}}$.
One might expect that even a small error in estimating $m_{\mathrm{test}}$ would invalidate predictions, but this is not necessarily the case. As shown in Section 4.2, if the correct representation is chosen, the lower bits maintain a strong periodic signal. Since the model can prepare multiple RNS representations (as demonstrated in step i), a sufficiently close estimate of $m_{\mathrm{test}}$ allows these lower bits to guide the model toward the correct representation. Further discussion is provided in Section H.4 and step iii.
Step iii: Once the necessary features are prepared, subsequent layers implement the rest of the algorithm. As in the FM case, we observe a ladder pattern in digit-wise accuracy for lower digits at early token positions. This pattern, which demonstrates the copying bias, is visible in Figure 9(b2).
More careful inspection of Figure 9(b2) shows the model copies the lowest 5 digits. Interestingly, given the estimated $m_{\mathrm{test}}=2033$ from step ii, the model effectively reduces it to $2033/2^{5}\approx 63.53$. Futhermore, since the model operates on integers, it could round it to $64$, yielding $64\cdot 2^{5}=2048=m_{\mathrm{test}}$. Thus, the model in principle, can predict higher digits accurately without the exact value of $m_{\mathrm{test}}$.
While some algorithmic transition from lower to higher digits clearly occurs, the exact mechanism remains unclear. The sharp transition in per-digit performance at the $2^{5}$ digit in Figure 9(b2) supports this argument. Nevertheless, we believe the model applies an algorithm similar to FM for higher bits, which we will elaborate on in Section H.5.